OpenStudy (anonymous):

How do I turn x^3+6x^2+13x+10=0 into a Quadratic formula (ax^2+bx+c=0) so I can find the solutions. The question says to Solve for the shown equation and I'm TOTALLY stuck...

ganeshie8 (ganeshie8):

you familiar wid below goodies ? 1) Descartes rule of signs 2) rational root test

ganeshie8 (ganeshie8):

and, 3) synthetic division

ganeshie8 (ganeshie8):

if so, its a piece of cake for u. simply find one zero using above. then apply quadratic formula for the depressed polynomial

OpenStudy (anonymous):

I don't know what Descartes rule of signs is OR the rational root test :'O I kind of get synthetic division. Care to explain <3?

ganeshie8 (ganeshie8):

Sure :) they're easy to *get*, actually synthetic division is hard to get, compared to descartes/rational roots

ganeshie8 (ganeshie8):

Descartes rule of signs :- 1) Count number of sign changes in f(x)

ganeshie8 (ganeshie8):

x^3+6x^2+13x+10=0 how many sign changes do u see ? all are positive, so 0 sign changes. that means there are NO POSIATIVE zeroes for this polynomial

ganeshie8 (ganeshie8):

So, there has to be atleast ONE negative zero for this cubic

ganeshie8 (ganeshie8):

we're done wid Descartes. now we know that, there wont be any positive roots, but, there has to be atleast ONE negative root.

ganeshie8 (ganeshie8):

you wid me still ?

OpenStudy (anonymous):

Yeah, so far :3

OpenStudy (anonymous):

Just taking notes as we go.

ganeshie8 (ganeshie8):

good, rational root theorem tells this :- ``` rational zeros for any polynomial will be of form : p/q p are factors of constant term q are factors of leading coefficient ```

ganeshie8 (ganeshie8):

x^3+6x^2+13x+10=0 ^ ^

ganeshie8 (ganeshie8):

for this, constant term = 10 leading coefficient = 1

ganeshie8 (ganeshie8):

Now tell me, wat are the factors of 10 ? wat are the factors of 1 ?

OpenStudy (anonymous):

For 10 its 1, 2, 5, and 10. 1 is just one right?

ganeshie8 (ganeshie8):

Awesome ! but u forgot negative factors

ganeshie8 (ganeshie8):

For 10 its 1, 2, 5, and 10 AND ALSO, -1, -2, -5 and -10

ganeshie8 (ganeshie8):

since Descartes already told us, there wont be any POSITIVE zeroes, dont wry about POSITIVE factors

ganeshie8 (ganeshie8):

lets test and see if -2 is a zero

ganeshie8 (ganeshie8):

use synthetic division

ganeshie8 (ganeshie8):

x^3+6x^2+13x+10=0 -2 | 1 6 13 10

ganeshie8 (ganeshie8):

can u divide, and see if it leaves 0 remainder ?

OpenStudy (anonymous):

I did all the math and got 0 as the answer to the synthetic division. That means 0 remainder right?

ganeshie8 (ganeshie8):

yup! and wats the depressed quadratic ?

ganeshie8 (ganeshie8):

also, that means -2 is a zero/solution of given cubic

ganeshie8 (ganeshie8):

-2 | 1 6 13 10 | -2 -8 -10 ------------------------ 1 4 5 0

ganeshie8 (ganeshie8):

so the depressed quadratic is x^2 + 4x + 5

ganeshie8 (ganeshie8):

you can use quadratic formula on this, and find other two zeroes/solutions

OpenStudy (anonymous):

Was doing the quadratic equation and got this on the top part of the division. How do I handle it?\[-4\pm \sqrt{-4}\]

ganeshie8 (ganeshie8):

pull out 4 from radical first

ganeshie8 (ganeshie8):

\(-4\pm \sqrt{-4}\) \(-4\pm 2\sqrt{-1}\)

ganeshie8 (ganeshie8):

\(\sqrt{-1} = i\)

ganeshie8 (ganeshie8):

\(-4\pm \sqrt{-4}\) \(-4\pm 2\sqrt{-1}\) \(-4\pm 2i\)

ganeshie8 (ganeshie8):

leave it like that. Done forget the denominator !

OpenStudy (anonymous):

So now I'm at \[\frac{4\pm2i }{ 8 }\] Would I be accurate in saying the answer/solution is \[\frac{ 1 }{ 2 }\pm \frac{ 1 }{ 4 }i\]

ganeshie8 (ganeshie8):

Nope. u should get : \(\large \frac{-4\pm 2i}{2} \) \(\large -2 \pm i\)